It is known that quantity demanded increases by two units for each $1 increases in price. At a price of $5, quantity demanded is ten units.
a. What will be the quantity demanded if price is zero
b. Write an equation for quantity demanded as a function of price
C. Write an equation that expresses price as a function of quantity
d. Write an equation for total revenue.
1
Expert's answer
2016-11-24T12:53:14-0500
The quantity demanded increases by two units for each $1 increases in price. At P = $5 Qd = 10 units.
a. If P = 0, then the quantity demanded is Qd = 2*0 = 0 units.
b. An equation for quantity demanded as a function of price is Qd = 2P.
C. An equation that expresses price as a function of quantity is P = 0.5*Q.
d. An equation for total revenue is TR = P*Q = 0.5Q*Q = 0.5*Q^2.